11•j=1(mod 4). Find j

1
Expert's answer
2022-03-21T03:57:04-0400

Let us solve "11\cdot j=1(\mod 4)." This congruence is equivalent to "(11-12)\cdot j=1(\mod 4)," and hence "- j=1(\mod 4)." We conclude that "j=-1(\mod 4)=3(\mod 4)." Therefore, "j=3+4t," where "t\in \Z."